OpenStudy (anonymous):

Find the domain of the vector function r(t)=< ln(9t),sqrt(t+3), 1/(sqrt(15-t))>

OpenStudy (zarkon):

find the domain of all 3 functions...take the intersection

OpenStudy (anonymous):

i got (-3,infinity)U(-infinity,15) but it says I'm wrong.

OpenStudy (zarkon):

(0,15)

OpenStudy (anonymous):

its still wrong =[

OpenStudy (zarkon):

is \[r(t)=\left<\ln(9t),\sqrt{t+3},\frac{1}{\sqrt{15-t}}\right>\]?

OpenStudy (anonymous):

yeah thats the problem

OpenStudy (zarkon):

then (0,15) is the answer

OpenStudy (anonymous):

o.o y don't you include the -3

OpenStudy (zarkon):

what is the value of the first component if you plug in -3

OpenStudy (anonymous):

ln(-27)

OpenStudy (zarkon):

correct...and you cant take the ln of a negative number ...unless you want a complex numbe

OpenStudy (anonymous):

ohhh thanks!!!